Grade 12Organic Chemistry

calculate the osmotic pressure of 5%solution of cane sugar (sucrose) at 1.5C {R=0.082 litre atm K-1mol-1}

MD IMRAN KHAN
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
moles sugar = 5 g / 342 g/mol=0.0146 
mass solution = 100 g 
assuming d = 1 g/mL 
volume solution = 100 mL = 0.100 L 
concentration sugar = 0.0146/0.100 =0.146 M 
osmotic pressure = 0.146 x 0.08206 x 274.5 K=3.297 atm